OpenStudy (anonymous):

I have a couple questions on a calculus worksheet I have. What is the position of S(x) of a moving object if its velocity "v" is defined as v(x)=9x^2-4x+2 and S(0)=11? Which of the following is the solution to the differential equation dy/dx=(x-5)/(y+1); y cannot equal -1 I need to be guided through, I don't just want the answer please. Thanks!

OpenStudy (phi):

integrate \( s(x) = \int v(x) \ dx \) for the second , you separate the variables, put y and dy on one side, and the x's and dx on the other

OpenStudy (anonymous):

I integrated and found the c to be 11, is that right? making the new equation \[3x^2-2x^2+2x+11\] I didn't know why else they would give me an initial condition. Why would I get them on two different sides? I thought the point was to have the dy/dx isolated?

OpenStudy (phi):

yes, when x is 0, your equation gives 11

OpenStudy (phi):

one technique to solve diff eq's is to separate the variables. i.e. treat dy and dx as variables that we can separate

OpenStudy (anonymous):

Okay, so the answer to the first question is just 3x^2-2x^2+2x+11? So I tried to do what you did but I ended up looking at my page for about five minutes...

OpenStudy (anonymous):

@phi

OpenStudy (phi):

dy/dx=(x-5)/(y+1); \[ \frac{dy}{dx} = \frac{(x-5)}{(y+1)} \\(y+1) \ dy = (x-5) \ dx \] integrate \[ \int (y+1) \ dy = \int (x-5) \ dx \\ \frac{y^2}{2} + y = \frac{x^2}{2} -5x + c \] which if we work at it, can be put in the standard form for the equation of a circle.

OpenStudy (anonymous):

Okay, I understood everything you did all the way up until you started talking about circles. How would I change it to a circle if I have a +c, do I just tag it on to the end?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

And shouldn't both antiderivatives have the +c?

OpenStudy (phi):

oh, it's a hyperbola As for the arbitrary constant, there are "properties" of arbitrary constants you might want to learn. For arbitrary a and b 2a --> a (the idea is 2a is still arbitrary... can be any value) a+ 3 --> a (adding a constant is still arbitrary) a+b= a --> (it is still some arbitrary number) thus if we multiply through by 2, we get 2c , but rename it as c (still arbitrary) \[ \frac{y^2}{2} + y = \frac{x^2}{2} -5x + c \\ y^2 +2y = x^2 -10x +c \] if we "complete the square" https://www.khanacademy.org/math/algebra/quadratics/completing_the_square/v/ex1-completing-the-square we can write this as \[ (y+1)^2 = (x-5)^2 + c \\ (y+1)^2-(x-5)^2 = c \] and that is the equation of a hyperbola with center (5,-1) It's properties will depend on the value of c, and we need some "initial condition" to determine c.

OpenStudy (phi):

Any questions ?

OpenStudy (anonymous):

One last question: A projectile is fired straight up from a platform 24 ft above the ground, with an initial velocity of 500ft/sec. Its downward acceleration is 32 ft/sec^2. What is the equation for the height of the projectile above the ground as a function of time "t" t=0 when the projectile is fired?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

You can use calculus/physics \[ V(t) = \int -32 \ dt \\ V(t)= -32 t+C \] at t=0, the initial velocity is 500 ft/sec, so V(t)= 500 - 32 t next \[ h(t) = \int V(t) \ dt = \int 500 - 32 t \ dt \\ h(t) = 500 t - 16t^2 + C \] at t= 0, the height is 24, so C=24, and \[ h(t)= 24 + 500 t - 16 t^2 \]

OpenStudy (anonymous):

Oh, I got all the way to the position function, but then I didn't know what to do. So the height is 24 at t=0, well that makes sense now. Thanks!